RC Meeting and Run Plan, Saturday, February 23

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav
Jpsi007.png

RC Meeting Information

No Meeting today

Useful Links

Current Status

  • Continued running J/Psi on kinematic setting 3 high-t
  • The target alarm situation where the heater/loss of cooling power goes to zero was occurring more frequently Friday evening.

Current Run Plan

Notices in Effect for Shift Crew

  • Shift Leader is responsible for 50k replays. Even if the TO is actually doing them.
  • Use the script do_coin_50k.sh to run the 50k replays. This will run both arm's first 50k replay. Then after ~500k events run a 50k replay for events 550k to 600k.

If you want to manually run each script, you can use a third argument to the standard 50k scripts which is the event number.


Current Issues

  • HMS dipole diff alarm. Current readback seems to have an issue. The NMR looks fine and the data shows no changes, so we are just keeping an eye on it for now.
  • Target rotation controller is turned off. Notify Steve Lassiter before any required rotations.

Planned Accesses to the Hall

Planned Accelerator Configuration Changes

  • Beam Studies on Monday (in preparation for Hall A PV experiments) require no beam for 8 hours.
  • We expect to get beam back around 4 pm (but hopefully sooner).

Special Requests

  • Latif will access the Hall on Monday (during beam studies) for GEM tests. This will require running a different DAQ configuration. This will take "a few hours".